Julia bought 3 more pencils than Jane,and Janet bought one third. As many as Jane.If each pencil cost $0.50,and they spent $15.50.Find the number of pencils Jane bought.

Answer by stanbon(75887)

Julia bought 3 more pencils than Jane,and Janet bought one third as many as Jane.If each pencil cost $0.50,and they spent $15.50.Find the number of pencils Jane bought.
----
Quantity Eq:: u = a+3
Quantity Eq:: n = (1/3)a
----
Value Eq:: 0.50(u + a + n) = 15.50
----------
Modify the value equation:
u + a + n = 31
---
Substitute for u and n and solve for "a":
a+3 + a + (1/3)a = 31
----
(7/3)a = 28
(1/3)a = 4
a = 12 (# Jane sold)
u = a+3 = 15 (# Julia sold)
n = (1/3)a = 4 (# Nancy sold)
-----------------------------
Cheers,
Stan H.
------------